OpenStudy (anonymous):

What is the slope and y-intercept of the graph of -x + 2y = 6?

OpenStudy (anonymous):

y intercept is 3 and slope is 1/1 or 1

OpenStudy (anonymous):

get correct answer by solving for y y = (x+6)/2 = (1/2) x + 3 so slope is (1/2) and y-intercept is 3